### v1

@@ 4.2.1 Surface rotation
In our sample, some stars (e.g. TIC 445256790) show regular sinusoidal modulations in their light curves that are different from the typical multi-periodic pulsational behaviour. The signal does not represent a beating pattern. Rather, the amplitude of the pulsations is constant but the light curves reveal an added sinusoidal component with a period of a few days. When comparing the typical longer variability to the above derived estimate of the rotation frequency, we find them to match. Hence, the longer-term variability is caused by rotational modulation, a feature commonly observed in high-resolution spectroscopy and high-precision photometry among the brightest $\beta$ Cep stars (e.g. $\lambda$ and $\kappa$ Sco Uytterhoeven et al., 2004, 2005; Handler & Schwarzenberg-Czerny, 2013).
@@
In order to select all possible $\beta$ Cep stars with rotational modulation, we searched for matching signals among the pre-whitened independent frequencies and the Gaia ESP-HS spectroscopic rotation rate. For each star with $f_{\mathrm{rot}}\sin i$, we select the closest signal and accept it as the (provisional) rotation frequency if it satisfies $2.50.9<f/f_{\mathrm{rot}}<$ 2.5, allowing for $0.41.1\lesssim\sin i<$ 0.4 to include the unknown projection effect of the rotational axis in the line-of-sight. Given the large uncertainty on the projected rotation rate the chosen range might exclude some true rotational signals but it ensures a clean selection. Overall, we find 41 % ($74/182$) of our sample to show some variability within the probed range.
@@
Most of the potential photometric rotational signals are close to the estimated $f_{\mathrm{rot}}\sin i$ indicating inclinations closer to the equator than to the pole of the rotation axis. This skewed distribution stems from two selection effects favouring $\sin i\sim 1$ above $\sin i\sim 0$. Firstly, the inclination angles of least cancellation of low-degree tesseral and sectoral modes deliver $\sin i\in[0.8,1.0]$ (see Appendix 1 in Aerts et al., 2010, for a definition and value of these viewing angles). Moreover, for the fastest rotators the pulsation modes tend to be sectoral modes concentrated to a band around the equator, maximising the visibility of such pulsations from the equatorial region (e.g. Reese et al., 2006). Secondly, stars with low $\sin i$ fall below the detection threshold in the Gaia $v\sin i$ measurements and hence do not occur in the sub-sample of stars with this quantity available.
For the majority of stars, we cannot see a rotational signal in the full light curve because the beating among the pulsation modes is much stronger than the potential rotational modulation. Mono-periodic stars are the only ones for which the rotational modulation can easily be seen from the light curve. The modulation is likely of weak magnetic origin through chemical spots. However, the typical fraction of magnetic massive stars is less than 10 % (e.g. Grunhut et al., 2017), which is lower than the fraction of detected surface rotation signals. Hence, either the fraction of magnetic $\beta$ Cep stars is larger than among non-pulsating massive stars (as suggested e.g. by Hubrig et al., 2009) or the detected signals are not only connected to magnetism.
@@ 4.2.2 Influence of internal rotation on pulsations
Figure 5 shows the main frequency observed in the TESS light curves against the estimated projected surface rotation frequency. We find a clear positive correlation between the two properties, as anticipated. We only have a (rather uncertain) direct estimate of the projected value of the surface rotation frequency in the line-of-sight, while the pulsation frequency’s shift is mainly determined by the rotation in the region of the star where the mode has most of its probing power. The low-order modes in $\beta$ Cep stars are dominant in the envelope and hence mainly probe the rotation in that region, denoted here as $\Omega_{\rm env}$. This, and the fact that the shift of any non-radial mode frequency in the line of sight is proportional to $m\Omega_{\rm env}$, implies that the main pulsation frequency $f_{1}$ is not tightly related to the projected surface rotation frequency, explaining the scatter in Fig. 5.
Figure 5: Main pulsation frequency $f_{1}$ against the projected surface rotation frequency $f_{\mathrm{rot}}\sin i$ of the 162 stars with available $v\sin i$ in the Gaia ESP-HS data. A correlation between both quantities is visible. Uncertainties in the pulsation frequency are typically within the symbol size.
@@ 5 Mode identification from TESS and Gaia data
Identifying the degree $l$ of at least one pulsation mode is required to reduce degeneracy and perform meaningful asteroseismic modelling of $\beta$ Cep stars (Hendriks & Aerts, 2019). In addition, it is also of importance to obtain an observational distribution of the mode degrees to advance our understanding of the mode selection mechanisms. Hence, we attempt to identify the pulsation modes of the Gaia sample to (1) learn about the distribution of pulsation modes in $\beta$ Cep stars from a large sample based on space photometry, (2) provide the input for detailed future modelling of the most promising sample stars, (3) open pathways towards ensemble asteroseismology of $\beta$ Cep pulsators observed across the entire sky from combined TESS and Gaia space data.
@@ 5.1 Multi-colour photometry
Traditionally in ground-based work, the pulsation modes of $\beta$ Cep stars have been identified with either multi-colour photometry or time-resolved spectroscopy (see Introduction, Sect. 1). In multi-colour photometry the amplitude ratios are calculated from the amplitudes of the pulsation modes as measured in different photometric filters. The bluest filter is typically used as the reference filter because of its higher SNR. In our work, the highest SNR is provided by the TESS filter. Yet we decided to use the bluest filter as reference band to enable comparison with earlier literature work on photometric mode identification (cf. Heynderickx et al., 1994). The lowest degree p-modes (i.e., radial modes with $l=0$) show a stronger drop-off with increasing wavelength than dipole or quadrupole modes, providing information about the mode degree $l$ (we refer to Heynderickx et al., 1994, for details of this methodology). As already discussed in Hey & Aerts (2024), the Gaia time series photometry offers multi-colour photometry which may help in mode identification.
@@
Figure 6 shows amplitude ratios of the Gaia $G_{\mathrm{RP}}$ and TESS $T$ band observations of stars with the same dominant pulsation frequency in Gaia and TESS. The ratios are computed with respect to the $G_{\mathrm{BP}}$ amplitude. In order to understand the position of different stars in this diagram and assess the feasibility of our approach, we marked stars with unambiguous known mode identifications from the literature ($l=1$: TIC 128821888, TIC 293680998, TIC 436285033; $l=2$: TIC 169786455, TIC 288489491, TIC 327656776). As expected the quadrupole ($l=2$) modes fall in the top right corner with amplitude ratios close to unity independent of the filter. Identified dipole ($l=1$) modes are found at lower amplitude ratios, while we expect to find the stars with dominant radial ($l=0$) mode in the lower left corner because their amplitudes decrease strongest with wavelength (see Heynderickx et al., 1994).
The few $\beta$ Cep stars with identified modes from long-term or multi-site campaigns in the literature tend to be either too bright to have Gaia time series photometry or have a low amplitude ($<4$ mmag). Hence no stars with known dominant radial pulsation is marked in Fig. 6. No clearly separated groups emerge from Fig. 6 in terms of the mode degree. Hence, we cannot achieve unambiguous identification from Gaia and TESS multi-colour amplitude ratios alone. However, the trend of having $l=2$ in the upper right corner and $l=1$ in the centre of the plot does emerge and is encouraging.
5.2 Search for rotationally split multiplets in the TESS data
Given the high photometric precision of the TESS light curves of these high amplitude $\beta$ Cep pulsators, mode identification might also be possible from the pre-whitened pulsation frequencies in the TESS data by identifying rotationally split multiplets, following the prototypical case of HD 129929 (Aerts et al., 2003). A pulsation mode with $l>0$ can be split into $12l+$ 1 observable mode components of azimuthal order $m$. In the corotating frame, the offset from the $m=0$ central mode is proportional to $\langle\Omega\rangle$ when treating the Coriolis force perturbatively up to first order (Ledoux, 1951).
Figure 7: Examples of rotationally split multiplets found in the amplitude spectra of the TESS data. The top panel shows an $l=1$ mode split into three azimuthal components. The central frequency of the multiplet ($f\approx 5.5$ d-1) had the dominant amplitude in the Gaia data. The strong peak left of the triplet corresponds is an independent mode. The bottom panel shows one of the few identified complete $l=2$ multiplets. In both panels the splittings are marked with grey markers connected by a line. The red markers at the top of each panel indicate the extracted frequencies from pre-whitening.
@@ 5.2 Search for rotationally split multiplets in the TESS data
Figure 7 shows examples of rotationally split multiplets we found in the TESS data. In total, we identified 12 incomplete and 48 complete rotational multiplets for the dominant frequency. The majority of these stars (83 %) pulsate in $l=1$ modes and far fewer in $l=2$ modes (17 %). However, we note that some of the potential modes are near the detection limit. Hence, only a subset of 37 of these stars with very clear splittings is used in the following analysis.
@@
Figure 8: Violin plots of amplitude ratios for identified pulsation modes over the filter’s effective wavelengths. The lines connect the mean values for each mode degree. Radial modes ($l=0$, green) are cleanly separated from non-radial pulsation modes ($l=1$, blue, $l=2$ orange), which have a larger overlap. The filter names are given at the top.
@@
Figure 8 shows the distributions of the mode degrees identified in this way over the effective wavelength of the considered filters [^2] All filters are broadband filters that overlap strongly. The overlap of the distributions are partly attributed to this overlap in wavelength.. This diagram includes only the stars for which we visually identified the pulsation modes with high certainty. For each photometric band and pulsation mode, we show the mean value of the amplitude ratio and a violin plot of the distribution. For all three considered degrees $l$ the distributions overlap strongly, which explains why we could not find a clean separation in Fig. 6 from the Gaia data alone and had to work with extra information from the rotational splitting in the TESS data. We find that the mean amplitude ratio values are distributed as expected, with a larger amplitude ratio for higher $l$. The radial modes show clearly smaller amplitude ratios than the non-radial modes. The distinction between $l=1$ and $l=2$ modes is small, following the results in Fig. 6.
@@ 5.3 Probabilistic mode identification from TESS and Gaia
With the secured mode identifications from rotational splittings in the TESS data, we now return to the Gaia-TESS multi-colour photometry for all pulsators, to identify the dominant mode of the stars without rotational splitting. We do so from their Gaia photometry following Fig. 6, after careful inspection of the position of the pulsators with rotational multiplets in this diagram.
@@
Given the overlap of the distributions of the amplitude ratios for modes of $l=0,1,2$ in Fig. 8 and the spread in Fig. 6, we aimed to estimate the mode identifications for the full sample based on a probabilistic framework including all photometric measurements and their uncertainties. To find the probability of a given mode identification under the measured Gaia and TESS amplitude ratios, we used the distributions shown in Fig. 8 as our training set and modelled them with a 3D Gaussian kernel density estimator ($K_{l}$). For each star, we calculated the likelihood $L_{l}$ for a given mode $l$ in a Naïve Bayesian scheme (Zhang, 2004) by integrating $K_{l}$ over the uncertainty interval of the measured amplitude ratios $\mathbf{A}$ of the three considered photometric bands (with $\sigma$ as their uncertainties):
$L_{l}=\int\int\int^{\mathbf{A}+\sigma}_{\mathbf{A}-\sigma}K_{l}(\mathbf{A})\mathrm{d}\mathbf{A}.$ (1)
The kernel density estimators were based on 9 values for $l=0$, 39 values for $l=1$, and 12 values for $l=2$.
To estimate the probability, we initially chose a flat prior but noticed that some stars with clear splittings were assigned $l=0$. Given the high occurrence rate of $l=1$ modes, we chose to use a non-flat prior with a weighting of 0.5 for $l=1$ and 0.25 for the other two considered degrees.
We identified 176 mode degrees ($l=0$: 44, $l=1$: 98, $l=2$: 34) with probability $p>50\,\%$. For 54 stars, we found a different mode degree to be more likely than the one assigned from visual inspection. Five of these stars were in our training set. Except for TIC 436240475, our initial identification had a relatively high probability close to 50 % as these stars are located in the overlap region of the distributions of two mode degrees. For reference and to inform future modelling efforts, we provide all mode identifications and their probabilities in Table 1.
As a validation of our methodology, we also reconsidered the stars with identified modes from the literature. All but one of these were assigned the correct mode degree. The exception, TIC 419354107 (BW Vul), is a very high-amplitude non-linear radial pulsator (Sterken & Jerzykiewicz, 1993) whose mode also got wrongly identified from high-resolution time-series spectroscopy when relying on linear pulsation theory (Aerts et al., 1995). Its position in Fig. 6 is indeed not consistent with the expectations for a linear radial mode, which is why we did not include it into our training set.
In summary, we are able to identify the dominant mode degree of 176 Gaia $\beta$ Cep pulsators with a probability $p>50\,\%$ based on their TESS light curve, Gaia multi-colour amplitude ratios, and the estimated surface rotation rates. This represents the largest sample of identified pulsation modes in $\beta$ Cep stars to date and can serve as a prime sample for further and detailed asteroseismic population studies of $\beta$ Cep stars. In the following, we present global seismic analyses and modelling of this homogeneously composed sample.
@@
TIC GaiaDR3 source ID RA Dec. $\log T_{\mathrm{eff}}$ $\log L$ Ref. $f_{\mathrm{rot}}\sin i$ $f_{1}$ $f_{\emph{Gaia}}$ $f_{\mathrm{TESS}}$ LC type $a_{G}/a_{\mathrm{BP}}$ $a_{\mathrm{RP}}/a_{\mathrm{BP}}$ $a_{T}/a_{\mathrm{BP}}$ $l$ $p(l=0)$ $p(l=1)$ $p(l=2)$ splitting $\delta_{A}$
(deg.) (deg.) (d-1) (d-1) (d-1) (d-1) (d-1)
4030832 5933776187446650880 246.18895 -53.46433 4.48 4.72 2 0.54 5.35610 5.35473 5.35610 mono 0.96 0.89 0.66 1 0.13 0.69 0.18 0.12 0.02
7038369 5933980456100089856 246.58808 -51.79895 4.35 3.93 0.19 4.46055 4.45906 4.46055 multi 0.91 0.77 0.57 1 0.30 0.67 0.03
11478661 2060836535411106304 303.93684 37.71048 4.32 4.07 0.21 6.28892 6.28626 6.28892 multi 0.57 0.64 0.33 0 1.00 0.00 0.00
11856739 5940654422886567936 247.12647 -49.35147 4.45 4.55 0.38 4.47288 4.47143 4.47288 hybrid 0.89 0.78 0.50 0 0.58 0.41 0.01
13332837 2061190956100233216 305.39636 38.61325 4.37 4.17 3, 4 0.48 6.87714 6.87450 6.87714 hybrid 0.90 0.73 0.65 1 0.13 0.83 0.04
13765605 465510069033121152 37.40739 61.20823 4.41 4.16 0.34 4.92968 4.92960 4.92968 multi 1.18 0.95 0.73 0.07 0.47 0.46 0.53 -0.10
13973539 2061131715608660736 305.76197 38.45576 4.40 4.65 3 0.26 3.50524 3.50277 3.50524 hybrid 1.03 1.01 0.63 1 0.15 0.54 0.31
14085632 2057943789022547968 305.72188 37.11278 4.53 4.94 0.38 4.63305 4.63295 4.63305 multi 0.94 0.93 0.89 2 0.03 0.03 0.94
15166556 5990434159009247232 244.59709 -45.84034 4.36 3.73 2 0.76 6.57384 6.57244 6.57384 multi 1.01 0.95 0.73 0.27 0.35 0.38 0.20 0.03

### v2

@@ 4.2.1 Surface rotation
Rotation leaves a signature in a star’s light curve by bringing chemical spots in and out of view. Some stars in our sample (e.g. TIC 445256790, HD 236535) show this kind of regular sinusoidal modulation in their light curves (Fig. 27) which is different from the typical multi-periodic pulsational behaviour. Rather, the amplitude of the pulsations is constant but the light curves reveal an added sinusoidal component with a period of a few days. When comparing the typical longer variability to the above derived estimate of the rotation frequency, we find them to match. Hence, the longer-term variability is caused by rotational modulation, a feature commonly observed in high-resolution spectroscopy and high-precision photometry among the brightest $\beta$ Cep stars (e.g. $\lambda$ and $\kappa$ Sco Uytterhoeven et al., 2004, 2005; Handler & Schwarzenberg-Czerny, 2013).
@@
In order to select all possible $\beta$ Cep stars with rotational modulation, we searched for matching signals among the pre-whitened independent frequencies and the Gaia ESP-HS spectroscopic rotation rate. For each star with $f_{\mathrm{rot}}\sin i$, we select the closest signal and accept it as the rotation frequency if it satisfies $2.50.9<f/f_{\mathrm{rot}}<$ 2.5, allowing for $0.41.1\lesssim\sin i<$ 0.4 to include the unknown projection effect of the rotational axis in the line-of-sight. Given the large uncertainty on the projected rotation rate the chosen range might exclude some true rotational signals but it ensures a clean selection. Overall, we find some variability within the probed range in 74 out of the 182 (41%) stars with an estimate for $f_{\mathrm{rot}}\sin i$.
@@
Most of the potential photometric rotational signals are close to the estimated $f_{\mathrm{rot}}\sin i$ indicating inclinations closer to edge-on. This skewed distribution stems from two selection effects favouring $\sin i\sim 1$ over $\sin i\sim 0$. Firstly, the inclination angles of least cancellation of low-degree tesseral and sectoral modes deliver $\sin i\in[0.8,1.0]$ (see Appendix 1 in Aerts et al., 2010, for a definition and value of these viewing angles). Moreover, for the fastest rotators the pulsation modes tend to be sectoral modes concentrated to a band around the equator, maximising the visibility of such pulsations from the equatorial region (e.g. Reese et al., 2006). Secondly, stars with low $\sin i$ fall below the detection threshold in the Gaia $v\sin i$ measurements and hence do not occur in the sub-sample of stars with this quantity available.
A substantial fraction (41 %) of our sub-sample features detectable rotational modulation, likely of magnetic origin through chemical spots. This is a higher fraction than the typical $\sim$10 % of magnetic field detections in spectro-polarimetric observations of massive stars (e.g. Grunhut et al., 2017). Hence, either the fraction of magnetic $\beta$ Cep stars is larger than among non-pulsating massive stars (as suggested e.g. by Hubrig et al., 2009) or space photometry is a more sensitive detection diagnostic or the detected signals are not only connected to magnetism.
@@ 4.2.2 Influence of internal rotation on pulsations
Figure 6 shows the main frequency observed in the TESS light curves against the estimated projected surface rotation frequency. We find a positive correlation between the two properties, as anticipated. We only have a (rather uncertain) direct estimate of the projected value of the surface rotation frequency in the line-of-sight, while the pulsation frequency’s shift is mainly determined by the rotation in the region of the star where the mode has most of its probing power. The low-order modes in $\beta$ Cep stars are dominant in the envelope and hence mainly probe the rotation in that region, denoted here as $\Omega_{\rm env}$. This, and the fact that the shift of any non-radial mode frequency in the line of sight is proportional to $m\Omega_{\rm env}$, implies that the main pulsation frequency $f_{1}$ is not tightly related to the projected surface rotation frequency, explaining the scatter in Fig. 6.
Figure 6: Main pulsation frequency $f_{1}$ against the projected surface rotation frequency $f_{\mathrm{rot}}\sin i$ of the 162 stars with available $v\sin i$ in the Gaia ESP-HS data. Uncertainties in the pulsation frequency are typically within the symbol size.
@@ 5 Mode identification from TESS and Gaia data
Identifying the degree $l$ of at least one pulsation mode is required to reduce degeneracy and perform meaningful asteroseismic modelling of $\beta$ Cep stars (Hendriks & Aerts, 2019). Additionally, an observational distribution of the mode degrees could advance our understanding of the mode selection mechanisms. Hence, we attempt to identify the pulsation modes of the Gaia sample to (1) learn about the distribution of pulsation modes in $\beta$ Cep stars from a large sample based on space photometry, (2) provide the input for detailed future modelling of the most promising sample stars, (3) open pathways towards ensemble asteroseismology of $\beta$ Cep pulsators observed across the entire sky from combined TESS and Gaia space data.
@@ 5.1 Multi-colour photometry
How a mode’s amplitude changes with wavelength depends on its degree, with low-degree modes showing a stronger decrease with wavelength. This has historically been used to identify pulsation modes of $\beta$ Cep stars using ground-based multi-colour photometry or time-resolved spectroscopy (see Introduction, Sect. 1). In multi-colour photometry the amplitude ratios are calculated from the amplitudes of the pulsation modes as measured in different photometric filters. The bluest filter is typically used as the reference filter because of its higher SNR. In our work, the highest SNR is provided by the TESS filter. Yet, we decided to use the bluest filter as reference band to enable comparison with earlier literature work on photometric mode identification (cf. Heynderickx et al., 1994). As already discussed in Hey & Aerts (2024), the Gaia time series photometry offers multi-colour photometry which may help in mode identification.
@@
Figure 7 shows amplitude ratios of the Gaia $G_{\mathrm{RP}}$ and TESS $T$ band observations of stars with the same dominant pulsation frequency in Gaia and TESS. The ratios are computed with respect to the $G_{\mathrm{BP}}$ amplitude. In order to understand the position of different stars in this diagram and assess the feasibility of our approach, we marked stars with unambiguous known mode identifications from the literature ($l=1$: TIC 128821888 (12 Lac), TIC 293680998 (IL Vel), TIC 436285033; $l=2$: TIC 169786455 (V856 Cen), TIC 288489491 (SY Equ), TIC 327656776 (KZ Mus)).Merely six stars in our sample have an identified mode from long-term or multi-site campaigns in the literature since these observations focussed on stars that are either too bright for Gaia or have an amplitude below $4$ mmag. Only TIC 419354107 (BW Vul) has an identified dominant radial pulsation in the literature, though its amplitude is so large that non-linear effects occur (Aerts et al., 1995), which may affect its amplitude ratios. Hence no stars with known dominant linear radial mode are marked in Fig. 7.
As expected the quadrupole ($l=2$) modes fall in the top right corner with amplitude ratios close to unity independent of the filter. Identified dipole ($l=1$) modes are found at lower amplitude ratios, while we expect to find the stars with dominant radial ($l=0$) mode in the lower left corner because their amplitudes decrease strongest with wavelength (see Heynderickx et al., 1994).No clearly separated groups emerge from Fig. 7 in terms of the mode degree. Hence, we cannot achieve unambiguous identification for any given star from Gaia and TESS multi-colour amplitude ratios alone. However, we also use the sample properties and additional information from the light curves. Particularly the trend of having $l=2$ in the upper right corner and $l=1$ in the centre of the plot does emerge and is encouraging.
5.2 Rotationally split multiplets and stillstand stars in the TESS data
Given the high photometric precision of the TESS light curves of these high-amplitude $\beta$ Cep pulsators, mode identification might also be possible by identifying rotationally split multiplets. Following the prototypical case of HD 129929 (Aerts et al., 2003), we use the pre-whitened pulsation frequencies in the TESS data. A pulsation mode with $l>0$ can be split into $12l+$ 1 observable mode components of azimuthal order $m$. In the corotating frame, the offset from the $m=0$ central mode is proportional to $\langle\Omega\rangle$ when treating the Coriolis force perturbatively up to first order (Ledoux, 1951).
Figure 8: Examples of rotationally split multiplets found in the amplitude spectra of the TESS data. The top panel shows an $l=1$ mode split into three azimuthal components. The bottom panel shows one of the few identified complete $l=2$ multiplets. The splittings are marked with grey markers connected by a line. The red markers at the top of each panel indicate the extracted frequencies from pre-whitening. The dashed lines mark the dominant frequency in the Gaia DR3 light curve.
@@ 5.2 Rotationally split multiplets and stillstand stars in the TESS data
Figure 8 shows two examples of rotationally split multiplets we found in the TESS data. For the example in the top panel, the dominant frequency in the Gaia DR3 light curve matches a significant frequency in the TESS light curve, where it is not the dominant one due to strong beating effects. In total, we identified 19 stars with a rotational multiplet involving the dominant Gaia frequency in the TESS data satisfying at least four of the following five criteria: 1) the multiplet is complete, 2) the splitting is on the same order of magnitude as $f_{\mathrm{rot}}\sin\,i$, 3) the splittings are symmetric up to $|\delta_{A}|<0.1$ (see Sect. 6.2), 4) the multiplet and its splittings do not conflict with any other potential multiplet in the Fourier spectrum, and 5) none of the frequencies in the multiplets originated from contaminating sources. The majority of these stars (17, representing 89 %) pulsate in a dominant $l=1$ mode and only two in a $l=2$ mode (representing 11 %). As we only seek three evenly spaced frequencies to identify an $l=1$ mode compared to five for an $l=2$ mode, the preponderance of $l=1$ modes may suffer from selection bias. In addition, we identified 14 stars with potential split multiplets that almost satisfied four conditions. We will come back to them after the mode identification in Sect. 5.3.
@@
Figure 9: Violin plots of amplitude ratios for identified pulsation modes over the filter’s effective wavelengths. The lines connect the mean values for each mode degree. The filter names are given at the top.
Unlike split multiplets for pulsators with $l\geq 1$, radial modes cannot easily be identified from the periodograms. However, the light curve of a high-amplitude radial pulsation in $\beta$ Cep stars sometimes exhibits the stillstand phenomenon, just as for Cepheids (e.g., Bersier et al., 1994). During the rise of the light curve, at the time of maximum radial velocity, the brightness briefly stalls before increasing further and dropping rapidly after the maximum (Walker, 1954). The stillstand phenomenon stems from the development of shock fronts in the expanding photosphere (Mathias et al., 1998; Fokin et al., 2004). The most prominent example among the known $\beta\,$Cep stars revealing a stillstand in its photometric light and spectroscopic radial-velocity curves is the high-amplitude non-linear radial pulsator BW Vul (TIC 419354107, e.g. Sterken et al., 1987; Aerts et al., 1995). Based on phase-folded and binned TESS light curves, we identified five more stars with stillstands (TIC 199208037, TIC 274070426, TIC 347520615, TIC 395218466, TIC 444357430). The folded TESS light curve of TIC 274070426 is shown alongside the one of BW Vul in Fig. 10, while those of the other four are displayed in Fig. 40. To our knowledge, none of these stars have been reported before to show this phenomenon. TIC 395218466 is of particular interest as it is a member of an eccentric eclipsing binary as detailed in Appendix E.2.
Figure 10: Light curves folded over the dominant frequency of two stars with clear stillstands in our sample (stalling of the light curve near phase 0.5). The folded light curve is binned into 35 bins, shown as black circles, to which we fitted a sinusoidal with a fixed frequency for reference (red). The top panel shows the archetypical stillstand star BW Vul (Walker, 1954).
@@
Three additional stars have been identified to have a radial mode due to the position of its frequency with respect to those of split multiplets. In total, we thus identified nine radial mode pulsators. Figure 9 shows the amplitude ratio distributions of all the mode degrees identified from rotational splitting or a stillstand in the TESS data, as well as from securely identified modes from the literature. We show the ratios as a function of the effective wavelength of the considered filters [^4] All filters are broadband filters that overlap strongly. The overlap of the distributions is partly caused by this overlap in wavelength.. For each photometric band and pulsation mode, we show the mean value of the amplitude ratio and a violin plot of the distribution. For all three considered degrees $l$ the distributions overlap strongly, which explains why we could not find a clean separation in Fig. 7 from the Gaia data alone and had to work with extra information from the rotational splitting or stillstand in the TESS data. We find that the mean amplitude ratio values are distributed as expected, with a larger amplitude ratio for higher $l$. The radial modes show clearly smaller amplitude ratios than the non-radial modes. The distinction between $l=1$ and $l=2$ modes is small, following the results in Fig. 7.
@@ 5.3 Probabilistic mode identification from TESS and Gaia
With the secured mode identifications from rotational splittings and stillstands in the TESS data, we now return to the Gaia-TESS multi-colour photometry for all pulsators, to identify the dominant mode of the stars without rotational splitting. Given the overlap of the distributions of the amplitude ratios for modes of $l=0,1,2$ in Fig. 9 and the spread in Fig. 7, we aimed to estimate the mode identifications for the full sample based on a probabilistic framework including all photometric measurements and their uncertainties. To find the probability of a given mode identification under the measured Gaia and TESS amplitude ratios, we used the distributions shown in Fig. 9 as our training set and modelled them with a 3D Gaussian kernel density estimator (KDE, $K_{l}$). As only very few dominant pulsation modes have been identified with a degree $l>2$ from ground-based observations (e.g. Heynderickx et al., 1994), we assumed $l\in\{0,1,2\}$. By cross-validation, we ensured that a Gaussian KDE indeed describes our data best. Similarly, we used cross-validation to choose the optimal value for KDE bandwidth parameter.
Since our training set is sparse, we augmented our observations by bootstrapping. In order to account for correlations in the data, we resampled the underlying pulsation amplitudes using a Gaussian perturbation based on the measurement uncertainties. We then calculated the amplitude ratios for each augmented data point. For every securely identified mode, we added 1 000 points to capture the uncertainty distribution properly. Based on the KDE calculated over the augmented data set and for every star, we calculated the likelihood $L_{l}$ for a given mode $l$ in a Naïve Bayesian scheme (Zhang, 2004) by integrating $K_{l}$ weighted by a Gaussian distribution $N$ with the uncertainties as its covariance $\mathbf{\sigma}$:
$L_{l}=\int\int\int K_{l}(\mathbf{A})N(\mathbf{A},\mathbf{\sigma})\mathrm{d}\mathbf{A}.$ (3)
To estimate the probability for the mode identification, we initially chose a flat prior but noticed that many stars with clear $l=1$ splittings were assigned either the correct degree but with a low probability or the wrong degree. This is an effect of the overlap of the distributions which reduces the probability of $l=1$ modes. Given the overall clear separation of $l=1$ modes, we chose to use a non-flat prior. The best prior is the one that enables the correct recovery for all known mode identifications. From a grid search of possible priors, we found the weights 0.33, 0.47, 0.20, for radial, dipole, and quadrupole mode occurrences, respectively, to provide the best result.
With the weighted probabilities, we identified 143 mode degrees ($l=0$: 23, $l=1$: 104, $l=2$: 16) with probability $p>60\,\%$. The large number of $l=1$ modes follows from the abundance of triplets in the training set and their prior probability. Five stars from the training set do not reach the chosen threshold of $p>0.6$ but only one of them is assigned the wrong degree. We have thus defined a robust probabilistic mode identification methodology suitable to treat all sample stars. For reference and to inform any future independent modelling efforts of the stars by community members, we provide all mode identifications (including those with $p<0.6$) and their probabilities in Table 1.
@@
TIC GaiaDR3 source ID Name RA Dec. G $\log T_{\mathrm{eff}}$ $\log L$ Ref. $f_{\mathrm{rot}}\sin i$ $f_{1}$ $f_{\emph{Gaia}}$ $f_{\mathrm{TESS}}$ LC type $P_{\mathrm{beat}}$ $A_{\mathrm{var}}$ $a_{G}/a_{\mathrm{BP}}$ $a_{\mathrm{RP}}/a_{\mathrm{BP}}$ $a_{T}/a_{\mathrm{BP}}$ $l$ $p(l=0)$ $p(l=1)$ $p(l=2)$ splitting $\delta_{A}$
(deg.) (deg.) (mag) (d-1) (d-1) (d-1) (d-1) (d) (d-1)
4030832 5933776187446650880 HD 147421 246.18895 -53.46433 9.02 4.48 4.72 2 0.54 5.35610 5.35473 5.35610 mono 0.21 0.98 0.96 0.89 0.66 1 0.09 0.73 0.18
7038369 5933980456100089856 246.58808 -51.79895 9.58 4.35 3.93 0.19 4.46055 4.45906 4.46055 multi 10.37 0.89 0.91 0.77 0.57 0.45 0.52 0.03
11478661 2060836535411106304 303.93684 37.71048 10.93 4.32 4.07 0.21 6.28892 6.28626 6.28892 multi 21.79 0.89 0.57 0.64 0.33 0 0.77 0.23 0.00
11856739 5940654422886567936 247.12647 -49.35147 10.78 4.45 4.55 0.38 4.47288 4.47143 4.47288 hybrid 0.37 0.91 0.89 0.78 0.50 0.52 0.48 0.00
13332837 2061190956100233216 HD 229085 305.39636 38.61325 9.74 4.37 4.17 3, 4 0.48 6.87714 6.87450 6.87714 hybrid 0.23 0.88 0.90 0.73 0.65 0.39 0.48 0.13
13765605 465510069033121152 37.40739 61.20823 11.15 4.41 4.16 0.34 4.92968 4.92960 4.92968 multi 2.10 0.94 1.18 0.95 0.73 1 0.01 0.66 0.34
13973539 2061131715608660736 HD 229171 305.76197 38.45576 9.18 4.40 4.65 3 0.26 3.50524 3.50277 3.50524 hybrid 0.75 0.90 1.03 1.01 0.63 1 0.08 0.75 0.17
14085632 2057943789022547968 305.72188 37.11278 10.67 4.53 4.94 0.38 4.63305 4.63295 4.63305 multi 7.21 0.88 0.94 0.93 0.89 2 0.02 0.18 0.80
15166556 5990434159009247232 HD 146442 244.59709 -45.84034 8.98 4.36 3.73 2 0.76 6.57384 6.57244 6.57384 multi 4.83 0.87 1.01 0.95 0.73 1 0.07 0.47 0.46 0.20 0.03
@@
Some stars in Fig. 7 exhibit a low $A_{T}/A_{\mathrm{BP}}$ amplitude, while near unity in $A_{\mathrm{RP}}/A_{\mathrm{BP}}$. A detailed analysis revealed that the majority of these stars contain a fraction of contaminating flux in their light curves. Although we tried to minimise the impact of contaminating light, it cannot always be avoided, for example in cases of increased sky background in TESS observations or very nearby bright stars as found in open clusters. Thanks to the two Gaia amplitude ratios included in the degree determination, the impact of contaminating light lowering the TESS amplitude is limited. Yet, it highlights the vulnerability of TESS amplitudes due to contamination from nearby sources and the sky background.
In summary, we are able to identify the dominant mode degree of 143 Gaia $\beta$ Cep pulsators with a probability $p>60\,\%$ based on their TESS light curve, Gaia multi-colour amplitude ratios, and the estimated surface rotation rates. For five additional stars, the multicolour photometry alone was inconclusive, yet we identified their mode degrees from splittings or a stillstand. This sample of 148 stars represents the largest sample of identified pulsation modes in $\beta$ Cep stars to date. Even if each of these pulsators revealed a limited number of identified pulsation modes compared to the handful of $\beta\,$Cep stars modelled in great detail (as summarised by Burssens et al., 2023), our results offer a prime sample for further and more detailed asteroseismic population studies of $\beta$ Cep stars. Only this type of sample studies can offer suitable input to calibrate stellar evolution theories for galactic massive field stars as a whole. In the following, we present our initial global seismic analysis and grid modelling of this unique and homogeneously composed sample.